OpenStudy (anonymous):

-2x+8y=11 solve for x

OpenStudy (anonymous):

Subtract the term with the y from both sides. Divide by the coefficient of x to get x alone.

OpenStudy (anonymous):

i got 2x=11-8y

OpenStudy (anonymous):

It should be -2x=11-8y. Don't miss that negative sign in front of the 2. :) Now divide both sides by -2 to get that x alone.

OpenStudy (anonymous):

x=11-6y

OpenStudy (anonymous):

i got x=11-6y

OpenStudy (anonymous):

No. You have to divide both terms on the right by -2. And make sure you're dividing, not subtracting.

OpenStudy (anonymous):

So now i got x=11-4y

OpenStudy (anonymous):

OHHHH you divide everything by -2

OpenStudy (anonymous):

You're not dividing both terms by -2.